### v1

# Question-Driven Analysis and Synthesis: Building Interpretable Thematic Trees with LLMs for Text Clustering and Controllable Generation

Abstract
Unsupervised analysis of text corpora is challenging, especially in data-scarce domains where traditional topic models struggle. While these models offer a solution, they typically describe clusters with lists of keywords that require significant manual effort to interpret and often lack semantic coherence. To address this critical interpretability gap, we introduce Recursive Thematic Partitioning (RTP), a novel framework that leverages Large Language Models (LLMs) to interactively build a binary tree. Each node in the tree is a natural language question that semantically partitions the data, resulting in a fully interpretable taxonomy where the logic of each cluster is explicit. Our experiments demonstrate that RTP’s question-driven hierarchy is more interpretable than the keyword-based topics from a strong baseline like BERTopic. Furthermore, we establish the quantitative utility of these clusters by showing they serve as powerful features in downstream classification tasks, particularly when the data’s underlying themes correlate with the task labels. RTP introduces a new paradigm for data exploration, shifting the focus from statistical pattern discovery to knowledge-driven thematic analysis. Furthermore, we demonstrate that the thematic paths from the RTP tree can serve as structured, controllable prompts for generative models. This transforms our analytical framework into a powerful tool for synthesis, enabling the consistent imitation of specific characteristics discovered in the source corpus.
Introduction
Unsupervised analysis of text corpora is a fundamental challenge in artificial intelligence, yet state-of-the-art methods face a critical trade-off between performance and interpretability. While modern topic models like BERTopic (Grootendorst 2022) can cluster documents semantically, they describe these clusters with lists of keywords that require significant post-hoc human effort to interpret and often lack coherence, especially in data-scarce domains. This interpretability gap motivates the central question of our work: can the structured reasoning capabilities of modern Large Language Models (LLMs) be employed to move beyond keyword-based topics and generate an intrinsically interpretable thematic model of a corpus? Furthermore, can such a model not only analyze content but also provide a structured foundation for controllably synthesizing new text?
To address these questions, we introduce Recursive Thematic Partitioning (RTP), a novel framework that reframes text analysis as an interactive dialogue with an LLM. RTP recursively builds a binary tree by prompting the model to generate a single, divisive natural language question at each node, such as “Is this review primarily about food quality or customer service?”. This process yields a fully interpretable hierarchy where each partition is explained by an explicit, semantic rule. Unlike existing methods, RTP produces a model that is transparent by design and whose thematic logic can be directly understood and validated by a human user.
We demonstrate through a comprehensive set of experiments that RTP provides a powerful new tool for both analysis and synthesis, as follows:
  1. 1.
    By comparing our method with a BERTopic baseline, we show that the thematic hierarchies generated by RTP are easier to interpret and more often linked to actionable information within the dataset.
  2. 2.
    We establish the quantitative utility of the generated clusters by showing they serve as effective features in downstream classification tasks.
  3. 3.
    We introduce and validate a novel application: using the paths of the tree as a structured “thematic signature” for Controllable Thematic Generation (CTG). Our results reveal that this technique provides unparalleled control over the semantic content of generated text, effectively solving a key challenge in controllable synthesis.
This work presents a unified framework that bridges the gap between understanding and creating, shifting from statistical pattern discovery to interactive, knowledge-driven text modeling.
Related Work
The Interpretability Challenge in Topic Modeling
Analysis of large textual corpora often relies on topic modeling to cluster documents by content (Vayansky and Kumar 2020). A canonical approach, Latent Dirichlet Allocation (LDA) (Blei, Ng, and Jordan 2003), models documents as bags-of-words and identifies topics based on word co-occurrence. While widely used, LDA’s performance degrades in data-scarce regimes, such as with small datasets or short texts.
To address these limitations, modern methods like BERTopic (Grootendorst 2022) use powerful pre-trained embeddings from models like BERT (Devlin et al. 2019). After clustering document embeddings, BERTopic uses TF-IDF on the resulting clusters to extract representative keywords.
Despite their different mechanisms, both classic and modern topic models share a critical limitation: their outputs require a post-hoc human interpretation of keyword lists. This process can be challenging because the generated keywords can be incoherent or meaningless for a specific application (Chang et al. 2009), and their interpretation can be inconsistent due to analysts’ expertise (Arnold et al. 2016) or hyperparameter choices (Gillings and Hardie 2022). Even when LLMs are used as external interpreters for these topics, their utility can be limited by a lack of domain-specific knowledge (Rijcken et al. 2023). While LLMs can also be prompted to extract topics directly (Pham et al. 2023; Mu et al. 2025), this often requires ad-hoc human intervention. This persistent interpretability gap motivates the central question of our work: can the structured reasoning capabilities of modern LLMs be applied to produce a thematically coherent and inherently interpretable representation of a corpus?
LLMs for Structured Reasoning and Analysis
While the extent to which LLMs possess emergent world models or true reasoning capabilities is debated (Mirzadeh et al. 2025; Shojaee et al. 2025; Abbe et al. 2024), their practical utility in structured tasks is well-established. Providing factual knowledge as part of the prompt, for example, can improve control and reduce hallucinations (Brown et al. 2020; Lewis et al. 2020). Also, explicitly inducing thought processes has shown to improve output quality and interpretability, as is the case of Chain-of-Thought (Wei et al. 2022) and Step-Back (Zheng et al. 2023) prompting.
A key capability that emerges from this grounded reasoning is the ability to answer specific questions about a text (Zhu et al. 2024) and organize these answers into structured formats. Prior work has established that this can be used to create interpretable classifiers from text (Balek et al. 2024). More relevant to our work, the ability of LLMs to generate structured outputs has been successfully applied to create binary decision trees from tabular data (Knauer et al. 2024; Xiong et al. 2024). This line of research is particularly promising, as tree-based models are not only highly interpretable but also consistently outperform other architectures on tabular data classification problems (Grinsztajn, Oyallon, and Varoquaux 2022).
Controllable Thematic Generation
A significant challenge in text generation is achieving fine-grained, reproducible control over content attributes. While LLMs can be prompted with adjectives like “formal” or “assertive”, these descriptions are often ambiguous and lead to inconsistent outputs (Schulhoff et al. 2024). Consequently, achieving high-fidelity imitation typically requires extensive prompt engineering or fine-tuning on large, specialized corpora (Liang et al. 2024).
The RTP Framework
Our work introduces Recursive Thematic Partitioning, a framework to build an interpretable binary tree from a corpus of unstructured text. The process, depicted in Figure 1, uses a greedy recursion similar to that of the well-known C4.5 decision tree (Quinlan 1993). It starts by associating all documents in the corpus to an initial, root node. Then, it recursively partitions a set of documents by generating a single, thematic question at each node and splitting the documents based on their answers. The implementation used in this work uses OpenAI’s gpt-4o-mini model, accessed via the Azure API.
Sample DatasetSummarize items (optional)Assign items to root nodeGenerate questionAnswer for Each ItemAssign Items to Child NodesRecurse into Child NodesLocal processLLM call
Figure 1: Overview of tree generation for RTP. A random sample of the dataset is taken. Then, the sample is used within a prompt to generate a yes/no question capable of partitioning the sample. Using LLM calls, the question is answered for each item in the sample. The items are assigned to the “yes” leaf and the “no” leaf, and the question generation process is recursively performed in each leaf until stop conditions are met.
Data Reduction
A primary practical challenge is that operating on a full corpus at each recursive step is bounded to the LLM’s context window size and to token costs. To create a tractable version of the dataset for the entire tree-building process, we employ a two-stage data reduction pipeline, which is executed before recursion begins:
  1. 1.
    Global Corpus Sampling. First, we draw a single, global random sample from the entire corpus. This smaller, sampled dataset is then used as the input for the complete recursive partitioning process. The sample size is a key hyperparameter governing the trade-off between representativeness and computational cost.
  2. 2.
    Document Summarization. Second, if the documents within our global sample are themselves very long, we can perform a one-time summarization step. Each document in the sample is summarized using an LLM call. We treat the use of summarization as a boolean hyperparameter.
This upfront data reduction strategy enables RTP to build a complete thematic hierarchy from a static, representative subset of the original large corpus.
Recursive Question Generation
At each node of the tree, starting with the root node containing the entire sampled dataset, we prompt an LLM to generate a single, divisive thematic question. The prompt contains a representative subset of documents from the current partition and instructs the LLM to create a yes/no question that best splits these documents into two coherent, roughly equal-sized groups.
Robust Data Partitioning
Once a question is generated, we prompt the LLM to generate the answer to the question for each document within the partition. To mitigate stochastic LLM errors, we employ a majority voting mechanism:
  1. 1.
    For each document, we query the LLM N times with a non-zero temperature, asking it to answer the generated question based on the document’s content.
  2. 2.
    The final answer (“yes” or “no”) for the document is determined by a simple majority vote over the N responses.
The number of votes, N, is a hyperparameter allowing a trade-off between robustness and token count. Documents are then split into a “yes” child node and a “no” child node based on their final assigned answer.
Recursion and Stopping Criteria
The process of question generation and partitioning is then recursively applied to both child nodes. The recursion for any given branch terminates when one of the following standard stopping criteria is met:
  1. 1.
    The tree reaches a predefined maximum depth, D.
  2. 2.
    The number of documents in a node falls below a minimum leaf size, K.
Both D and K are hyperparameters of the framework.
Classification
The generated tree can be used for classification. For such, the tree is used as a decision path. For a given document, we traverse the tree from the root, using an LLM to answer the question at each node to determine the path. The final prediction is the majority class of the training samples that landed in the corresponding leaf node.
Controllable Thematic Generation
Controllable Thematic Generation (CTG) leverages the explicit structure of the RTP tree to generate texts that adhere to specific, user-defined themes. The process begins by selecting a target leaf node that represents the desired thematic content. We then extract the “thematic signature” of this node, defined as the unique sequence of questions and answers along the path from the root. This sequence is then translated into a set of structured constraints that are incorporated into a prompt to guide a generative LLM.
Experiments and Results
Our experiments are designed to answer three central research questions regarding RTP:
  1. 1.
    Interpretability: Does RTP produce thematic hierarchies that are demonstrably more interpretable to humans than those from state-of-the-art topic models?
  2. 2.
    Quantitative Utility: Do the clusters generated by RTP serve as useful features in downstream supervised tasks?
  3. 3.
    Controllable Thematic Synthesis: Can the thematic paths of an RTP tree be used as structured prompts to controllably generate text within a specific, data-driven theme?
In the following sections, we describe the methods used to address each of these questions, followed by a detailed analysis of our findings.
Datasets
We evaluate our framework, RTP, on three widely-used public benchmarks: IMDB Movie Reviews  [^1] https://huggingface.co/datasets/stanfordnlp/imdb (Maas et al. 2011), Yelp Reviews  [^2] https://huggingface.co/datasets/Yelp/yelp˙review˙full (Zhang, Zhao, and LeCun 2015), and AG News  [^3] https://huggingface.co/datasets/fancyzhx/ag˙news (Zhang, Zhao, and LeCun 2015). These datasets were strategically chosen for their well-understood semantic content, which allows for a straightforward assessment of the interpretability and alignment of the generated themes. Moreover, the clear nature of sentiment in a movie review or the category of a news article enables a critical assessment of the results without requiring specialized domain knowledge.
Importantly, some items in each dataset triggered the LLM’s safety protections due to violent or sexual content. These items were discarded from the sample.
Qualitative Analysis of Thematic Trees
We generated thematic trees for each dataset to perform a qualitative analysis of RTP’s behavior [^4] Parameters: sample size 100, $N=4$, $D=5$, $K=3$, no summarization. They are shown in figures 2 (IMDB), 3 (Yelp Reviews), and 4 (AG-News). We immediately note that the questions seem to be related to the dataset content.
[root] Does the text primarily provide a positive review of the film or work discussed?
      [yes] Does the text describe the production or plot of a movie, show, or documentary that involves or depicts a romantic relationship between characters?
      [yes] Does this text describe a film where the main character faces a personal moral or ethical struggle?
      [no] Does the text describe a romantic relationship as a central theme in the movie?
      [yes] Does the text mention a prominent female character in the story?
      [no] Does the text discuss the cinematography or visual presentation of the film or show?
      [yes] Does the text mention a documentary or a film about saving the planet or environmental issues?
      [no] Does the text describe the film in a very positive light overall?
      [no] Does the text describe a film or show as being emotionally impactful?
      [yes] Does the text recommend watching the movie or show?
      [no] Does the text mention a television show or series?
      [no] Does the text describe the movie as disappointing or poorly executed?
      [yes] Does the reviewer criticize the script for being poorly written or uninteresting?
      [yes] Does the text have a positive overall assessment of the movie?
      [no] Does the text have a positive overall assessment of the movie?
      [no] Does the text mention disappointment with the film’s acting or storyline?
      [yes] Does the text criticize the film’s lack of coherence or clear direction?
      [no] Does the text describe the movie as boring?
      [no] Does the text mention a train?
      [no] Does the text mention a specific film director by name?
      [no] Does the text critique an unjust or exaggerated portrayal of characters or events?
Figure 2: Thematic tree generated for the IMDB dataset.
[root] Does the text mention the author having a positive experience with the food or services provided?
      [yes] Does the review mention being satisfied with the food at the establishment?
      [yes] Does the text mention a specific dish or food item that the author enjoyed at the establishment?
      [yes] Does the text mention a specific dish or menu item that the author enjoyed?
      [yes] Does the text mention a specific dish or menu item that the author enjoyed?
      [no] Does the text mention positive aspects of the customer service or staff?
      [yes] Does the reviewer mention exceptional or notable service in the text?
      [no] Does the text mention a specific issue or drawback related to the described experience or place?
      [no] Does the text mention dissatisfaction with food or drink quality?
      [yes] Does the review mention an issue with the food’s preparation or quality?
      [yes] Does the review mention an issue with the food’s preparation or quality?
      [no] Does the text describe a place where a purchase can be made?
      [yes] Does the text mention a specific price for a product or service?
      [no] Does the text mention experiencing poor customer service?
      [yes] Does the text describe an experience with poor service?
      [yes] Does the text describe an experience with poor service?
      [yes] Does the text describe an experience with poor service?
      [no] Does the text describe a positive experience with the food at the establishment?
      [no] Is this review primarily focused on dissatisfaction with the food or beverage quality?
      [yes] Does the text describe significant dissatisfaction with the food quality?
      [no] Does the text describe an experience involving food or drink?
Figure 3: Thematic tree generated for the Yelp Reviews dataset.
[root] Does the text describe a significant organizational, economic, or structural change?
      [yes] Does the text mention a plan or agreement to cut jobs, operations, or resources?
      [no] Does the text mention financial losses or challenges faced by a company?
      [yes] Does the text mention financial troubles or losses for the company involved?
      [yes] Does the text report a resolution or agreement in response to a financial or operational challenge?
      [no] Does the text mention a company cutting, reducing, or altering any of its offerings or practices?
      [yes] Does the text mention a reduction or change being made to a product or practice?
      [no] Does the text mention a merger or collaboration between two entities?
      [no] Does the text mention sports or athletes?
      [yes] Does the text mention a victory or winning event?
      [yes] Does the text mention a specific individual achieving victory or recognition?
      [yes] Does the text mention a victory or win in a sports competition?
      [no] Does the text describe a decisive event or moment in a sports context?
      [yes] Does the text specifically mention a score or ranking?
      [no] Does the text mention an injury?
      [no] Does the text explicitly mention a financial market or economic activity?
      [yes] Does the text mention oil prices or crude oil?
      [yes] Does the text mention oil prices changing?
      [no] Does the text mention a financial or economic statistic or figure?
      [no] Does the text mention a government action or decision?
      [yes] Does the text mention an instance of a government or official body negotiating or responding to a conflict or dispute?
      [no] Does the text mention a specific event occurring in nature or the environment?
Figure 4: Thematic tree generated for the AG-News dataset.
A qualitative examination of the trees generated by RTP reveals several key characteristics of our method, highlighting both its strengths in thematic discovery and its interesting emergent behaviors.
Hierarchical Theme Discovery
For the sentiment-focused IMDB and Yelp Reviews datasets, RTP consistently identifies the primary axis of variation. The root node in both trees is a question directly related to positive versus negative sentiment, confirming that RTP correctly captures the most dominant theme. Subsequent nodes in the hierarchy then explore more nuanced, fine-grained aspects within each sentiment class. For example, as shown in Figure 2, positive reviews in the IMDB dataset are further partitioned by questions about the movie’s emotional impact, while negative reviews are split by criticisms of technical aspects like cinematography or acting. Similarly, the Yelp tree reveals that negative customer experiences are often thematically clustered around poor customer service, which is a highly specific and actionable insight for a business owner.
In contrast, the root node of the AG News tree asks a question that separates articles related to geopolitics and finance from those concerning sports and technology. This demonstrates that RTP identifies semantic themes that may not perfectly align with the predefined ground-truth labels but still represent a valid and logical partition of the content. The resulting hierarchy thus provides a more fine-grained thematic view than the four broad categories alone.
Emergent Behaviors and Limitations
Our analysis also highlights several emergent behaviors that characterize the current RTP framework.
Unbalanced Partitions. While we prompt the LLM to generate questions that create roughly equal splits, we observed several nodes, particularly in the IMDB and Yelp trees, that result in highly unbalanced branches. This occurs when the LLM generates a very specific question that isolates a small but coherent thematic minority (e.g., reviews focused on a niche technical flaw). This indicates a failure to strictly adhere to the ”half-and-half” constraint, but it successfully identifies a distinct, interesting subgroup. This suggests a trade-off between creating balanced trees for exploration and isolating small, highly specific themes.
Thematic Redundancy. We also noted instances of similar or repeated questions appearing in the tree, either in parent-child relationships or between sibling nodes. This is a natural consequence of the algorithm’s design, where question generation for each node is an independent process that only considers the local document partition. The model has no memory of the questions asked previously in the hierarchy. While this can lead to redundancy, it also demonstrates the robustness of the LLM in identifying the most salient thematic split at different levels of granularity.
Summary of Analysis
Overall, our qualitative analysis confirms that RTP successfully builds a meaningful and interpretable thematic hierarchy. It correctly identifies the dominant themes at the root and uncovers nuanced, actionable sub-themes in deeper nodes. Its emergent behaviors, such as unbalanced splits and thematic redundancy, do not detract from its utility but rather provide valuable insights into the method’s decision-making process and highlight clear avenues for future work, such as incorporating tree-level context into the question generation prompt.
Interpretability: RTP vs. BERTopic
Setup and results
To provide a fair comparison with a standard topic model, we executed BERTopic in the same dataset sample used in the experiments with RTP. The results are shown in Tables 1, 2, and 3. Following BERTopic convention, topics are numbered from 0 to the number of topics found, and topic 1-1 represents outliers.
Topic Representation
-1 br movie like just really film best actors time joseph
0 br film movie like just bad good really movies thing
1 film br movie police hes mukhsin tolstoy way good love
2 film musical best brando singing dancing george burns astaire fontaine
3 br action point dont sasha film does believe like money
4 ethel movie dead story advancement backwoods bodies battalion halfway film
5 color life natali films think movie important based white kind
Table 1: Topics generated by BERTopic in IMDB
Topic Representation
-1 food like really just good place groupon ordered lunch told
0 bar got airport way people air like line view terminal
1 great good delicious ordered place places crepes breakfast lunch phoenix
2 food thai good great pho place menu hair love felt
3 room told hotel desk card security stay roland didnt safe
4 table horrible minutes busy servers sat right server service del
5 pizzas pizza papa called crust toppings johns driver looked time
6 wash sushi mat creme includes vacuum price machine priced tea
7 salmon hype fruit place garlic dont flavors instead fries time
8 salad cardboard hickory food salads piece know dressing husband came
9 buffet wait quality seated relleno chile youre enchiladas beans food
10 sandwiches night cute owner nni coffee lots quite huge big
Table 2: Topics generated by BERTopic in Yelp Reviews
Topic Representation
-1 ap monday bush biggest said solaris group british hp 39s
0 39s reuters oil corp prices cut stocks said new york
1 election iran nuclear said afp north ltbgtltbgt syria sudan region
2 lions 39t inter williams win sunday season standings dallas revolution
3 yankees sox look jeter games johnson peskys rays aftermath charleys
4 39s england world bowler open number test coach champions harmison
5 bison bears falls taj river salmon mahal mcneil spectacular sun
6 alitalia alter emissions end split practices canada unions deal sales
Table 3: Topics generated by BERTopic in AG-News
Analysis
We note that the keywords seem to be related to the dataset ideas. However, they do not convey as much information as the themes generated by RTP. For example, theme 0 in the IMDB dataset is related to both “good” and “bad”, whereas we can find topics that cite specific actors or actresses. Similarly, in the Yelp Reviews dataset there are mentions of specific dishes, and the topics for AG-News mention specific companies, cities, and sports teams. This showcases that RTP yields more general themes, while this generalization must be performed manually when using BERTopic.
However, it is important to note that BERTopic topics represent the result of an objective analysis of an embedding-based clustering of the collection, whereas the themes for RTP rely on a strong pre-training of a stochastic generative question-answering system. This means that each word in the topics created by BERTopic can be explained in terms of their role, whereas the themes created by RTP are more strongly linked to the datasets used to train the LLM itself. Consequently, although RTP-generated themes are easier to interpret, they can be misleading due to biases that cannot be controlled by the user.
Text Classification
Setup
To evaluate the quantitative utility of the thematic hierarchies generated by RTP, we use them as classifiers for the ground-truth labels of each dataset. Also, we further analyze the effects of summarization by turning it on and off both in the training and testing stages. The process is repeated five times in different samples of the training set and mean and the two-sigma confidence interval are reported.
To obtain a baseline result, we fine-tune a pre-trained DistilBERT model (Sanh et al. 2019) on the exact same small data sample used to generate the RTP tree. This ensures a fair comparison, isolating the benefit of RTP’s structure against the fine-tuning of a powerful neural network on the same limited data. For context, we also report the state-of-the-art (SoTA) performance from models on huggingface.com  [^5] IMDB: https://huggingface.co/DeepAxion/distilbert-imdb-sentiment; Yelp Reviews: https://huggingface.co/BexRedpill/distilbert-on-yelp-reviews-full-epoch-2; AG-News: https://huggingface.co/JiaqiLee/bert-agnews, which are typically trained on the full dataset.
Results
Our primary finding is that RTP significantly outperforms the fine-tuned DistilBERT baseline in IMDB, and surpasses the performance of a model trained on the full dataset ($p<0.05$, one-sample t-test). The performances of RTP and the baseline are similar in Yelp Reviews. However, on the topic-based AG News dataset, the fine-tuned DistilBERT baseline is superior. This brings forward a key characteristic of our method.
Model IMDB Yelp Reviews AG-News
SoTA (reference) .94 .65 .94
DistillBERT (baseline) .81(.03) .39(.07) 0.87(.01)
RTP (ours) 96(.02) .40(.12) .64(.04)
RTP-T (ours) .96(.02) .39(.04) .52(.10)
RTP-E (ours) .91(.03) .37(.08) .61(.06)
RTP-TE (ours) .89(.04) .41(.05)* .51(.07)
Table 4: Accuracies for each model. SoTA is the highest accuracy reported in models found in huggingface.com. The RTP classifiers are identified on whether they use summarization on training (T) or on evaluation (E). The highest mean accuracy (excluding SoTA) is marked in bold. Values are $\mu(2\sigma)$. Asterisks denote non-significant differences ($p>.05$) compared to baseline.
Analysis
The performance increase in IMDB can be attributed to the alignment between the question in the root node (“Does the text primarily provide a positive review of the film or work discussed?”) and the classification problem itself. Thus, the high classification accuracy is a consequence of the LLM’s ability to answer that question correctly. With an adequate set of questions, the classification process exploits the zero-shot learning ability of LLMs and can reach close to SoTA accuracies.
The root question in the Yelp Reviews tree (“Does the text mention the author having a positive experience with the food or services provided?”) is ambiguous with regards to the classification problem: a review can be primarily positive without explicitly mentioning the author’s experience. In AG-News, the questions throughout the tree are only indirectly aligned to the classification problem. Henceforth, the high classification accuracy is a consequence of the dataset partitions identified by RTP being correlated to the ground truth labels.
Our ablation study on summarization revealed that it has a different effect in each dataset. The use of summarization was harmful in IMDB and AG-News, whereas it had no statistically significant effect in Yelp Reviews. Also, summarization in training had a smaller impact in IMDB, while summarization on evaluation had a smaller impact on AG-News. Henceforth, the effects of summarization depend on the datasets themselves, and should be treated as a dataset-dependent hyperparameter.
Controllable Thematic Generation
Setup
The proposed Controllable Thematic Generation (CTG) was evaluated using the trees generated for IMDB, Yelp Reviews, and AG-News. We used two baselines. The first, a generic prompt that uses no information from the dataset, which we called “uncontrolled”. Second, we used a strategy based on few-shot prompting in which all texts in the target node were used as examples (instead of the explicit questions and answers that compose the theme signature). The generation process was executed for an arbitrary leaf node in each dataset, and evaluated using two strategies.
Results
The first evaluation strategy was to use Sentence Bert (Reimers and Gurevych 2019) to calculate the embedding for each document in the original leaf and in the generated samples [^6] We used the all-MiniLM-L6-v2 model. It operated in the initial tokens of each text and discarded those that did not fit the context window, under the assumption that style and tone are constant throughout each text. Then, the cosine similarity between the centroids of each generated sample and the original sample was calculated. The results, shown in Table 5, indicate a general coherence of writing style and phrasal similarity. Clearly, in this aspect, few-shot prompting is a far superior technique.
Strategy IMDB Yelp Reviews AG-News
Uncontrolled .37 .51 -.01
Few-shot (baseline) .45 .55 .86
CTG (ours) .41 .52 .34
Table 5: Sentence Bert embedding similarity evaluation of each generation strategy.
The second evaluation strategy was to use the classification algorithm to associate each document with a leaf node. A document associated with the same node used to start the generation was considered correct. This allows calculating an accuracy, as shown in Table 6, which shows that CTG outperforms other techniques.
Strategy IMDB Yelp Reviews AG-News
Uncontrolled .07 .00 .00
Few-shot (baseline) .04 .03 .08
CTG (ours) .60 .39 .31
Table 6: Node accuracy evaluation of each generation strategy.
Analysis
These contrasting results reveal an important finding. The choice of generation strategy depends on the user’s goal. Few-shot prompting is a powerful technique for imitating surface-level features like phrasing and tone. However, our CTG method provides control over thematic content, making it the superior tool when the goal is to generate text that adheres to specific, rule-based semantic constraints. This is because in few-shot prompting the LLM has the burden of extracting patterns from the dataset, whereas the CTG method specifically informs what are these patterns. This distinction between mimicking style and controlling substance is a key contribution of our work.
Conclusion
In this paper, we introduced Recursive Thematic Partitioning (RTP), a novel framework that uses Large Language Models to build interpretable thematic hierarchies from text corpora. By recursively partitioning data with natural language questions, RTP moves beyond the keyword-based outputs of traditional topic models, creating a transparent taxonomy where each decision point is explained. We demonstrated that this approach yields hierarchies that are not only more interpretable to humans but also quantitatively useful as features in downstream tasks. Furthermore, we showed that the analytical structure of the RTP tree can be repurposed for synthesis, providing a powerful, data-driven method for Controllable Thematic Generation that offers unparalleled control over content.
RTP opens exciting avenues for future research. A key next step is to investigate optimal strategies for integrating human-in-the-loop feedback, especially in niche contexts such as education and law. The explicit, question-based nature of each node creates natural intervention points where a domain expert could validate, override, or refine the LLM’s suggested partitions. Exploring interactive interfaces that allow users to guide the tree-building process could further enhance the model’s alignment with specific analytical goals, bridging the gap between automated discovery and expert-driven knowledge construction.

### v2

@@
 Each node in the tree is a natural language question that semantically partitions the data, resulting in a fully interpretable taxonomy where the logic of each cluster is explicit.
-Our experiments demonstrate that RTP’s question-driven hierarchy is more interpretable than the keyword-based topics from a strong baseline like BERTopic.
-Furthermore, we establish the quantitative utility of these clusters by showing they serve as powerful features in downstream classification tasks, particularly when the data’s underlying themes correlate with the task labels.
+Our experiments suggest that RTP’s question-driven hierarchy is more interpretable than the keyword-based topics from a baseline like BERTopic.
+Furthermore, we show that formed clusters can serve as powerful features in downstream classification tasks, particularly when the data’s underlying themes correlate with the task labels.
 RTP introduces a new paradigm for data exploration, shifting the focus from statistical pattern discovery to knowledge-driven thematic analysis.
@@
   1. 1.
-    By comparing our method with a BERTopic baseline, we show that the thematic hierarchies generated by RTP are easier to interpret and more often linked to actionable information within the dataset.
+    We demonstrate that RTP generates thematic hierarchies based on explicit, natural language questions.
+By contrasting our method with a keyword-based baseline like BERTopic, we show that this question-driven approach offers a transparent, rule-based form of interpretability where the logic of each partition is human-readable by design.
   2. 2.
-    We establish the quantitative utility of the generated clusters by showing they serve as effective features in downstream classification tasks.
+    We investigate the quantitative utility of the generated clusters as features for downstream classification.
+Our preliminary analysis highlights their potential in settings where the discovered themes align with the task’s underlying labels.
   3. 3.
-    We introduce and validate a novel application: using the paths of the tree as a structured “thematic signature” for Controllable Thematic Generation (CTG).
-Our results reveal that this technique provides unparalleled control over the semantic content of generated text, effectively solving a key challenge in controllable synthesis.
+    We introduce a novel application for the resulting thematic paths: Controllable Thematic Generation (CTG).
+We demonstrate how these “thematic signatures” can be used as structured prompts to exert fine-grained, rule-based control over the semantic content of generated text, distinguishing this approach from style-imitation techniques.
 This work presents a unified framework that bridges the gap between understanding and creating, shifting from statistical pattern discovery to interactive, knowledge-driven text modeling.
@@
 This smaller, sampled dataset is then used as the input for the complete recursive partitioning process.
-The sample size is a key hyperparameter governing the trade-off between representativeness and computational cost.
+The sample size $S$ is a key hyperparameter governing the trade-off between representativeness and computational cost.
+Also,
   2. 2.
@@
 Second, if the documents within our global sample are themselves very long, we can perform a one-time summarization step.
-Each document in the sample is summarized using an LLM call.
+Each document in the sample is summarized using an LLM call, using the prompt in Figure 2 as system instructions and the text to be summarized as user instructions.
 We treat the use of summarization as a boolean hyperparameter.
+System Prompt Summarize the following text into summary with a maximum of {max_words} words.
+The summary should be concise and capture the main points of the text.
+Reply solely with the summary, without any additional text or formatting.
+Your answer should be a JSON object following the schema: {schema_str} User Prompt Text: {text}
+Figure 2: Prompts used for summarization.
+The text data is provided to the LLM as a user prompt.
+By changing the max_words parameter, we can control the summary length.
+The JSON schema used for prompting helps parsing the answer.
 This upfront data reduction strategy enables RTP to build a complete thematic hierarchy from a static, representative subset of the original large corpus.
 Recursive Question Generation
-At each node of the tree, starting with the root node containing the entire sampled dataset, we prompt an LLM to generate a single, divisive thematic question.
-The prompt contains a representative subset of documents from the current partition and instructs the LLM to create a yes/no question that best splits these documents into two coherent, roughly equal-sized groups.
+At each node of the tree, starting with the root node containing the entire sampled dataset, we prompt an LLM to generate a single, divisive thematic question, using the system instructions shown in Figure 3.
+The user prompt contains a representative subset of documents from the current partition and instructs the LLM to create a yes/no question that best splits these documents into two coherent, roughly equal-sized groups.
+System Prompt You are an analyst.
+You will receive a list of texts formated as a list.
+You will propose one yes/no question whose answer can be determined from each individual text.
+The question must be answerable “yes” to half of these texts, and “no” to the other half.
+The question must be concise and clear, allowing for a straightforward yes/no answer for each text.
+Do not use any external knowledge or assumptions.
+Your answer must be determined only from the texts provided.
+You must reply following a json in the schema: {schema_str} User Prompt Texts: {texts}
+Figure 3: Prompts used for question generation.
+The text data is provided to the LLM as a user prompt containing a Python-style list of texts.
+The JSON schema used for prompting helps parsing the answer.
 Robust Data Partitioning
-Once a question is generated, we prompt the LLM to generate the answer to the question for each document within the partition. To mitigate stochastic LLM errors, we employ a majority voting mechanism:
+Once a question is generated, we prompt the LLM to generate the answer to the question for each document within the partition, using the prompt shown in Figure 4.
+System Prompt You will receive a text and a question about it.
+Answer the question as yes (true) or no (false) based solely on the text.
+Do NOT use any external knowledge or assumptions.
+Your answer must be determined only from the text provided.
+If the question is not answerable from the text, respond with ”false”.
+Ignore assumptions or implications not explicitly stated in the text.
+Reply with a JSON object following the schema: {schema_str} User Prompt Text: {text} Question: {question}
+Figure 4: Prompt used for data partitioning. The text data is provided to the LLM as a user prompt containing a Python-style list of texts. The JSON schema used for prompting helps parsing the answer.
+To mitigate stochastic LLM errors, we employ a majority voting mechanism:
   1. 1.
@@
 The number of votes, N, is a hyperparameter allowing a trade-off between robustness and token count.
+This parameter is more relevant when the questions and related documents are more ambiguous (for example: asking whether a news article narrating a football game is about sports is straightforward, whereas asking whether that same news article narrates overall positive news can be different depending on the reader’s opinion about the winning team).
 Documents are then split into a “yes” child node and a “no” child node based on their final assigned answer.
+Draws are arbitrarily assumed to be a “yes”.
+This could lead to biasing towards positive answers.
+However, draws tend to only occur if questions are ambiguous in relation to the analyzed text, which means either answer would be equally suitable.
 Recursion and Stopping Criteria
@@
 Both $D$ and $K$ are hyperparameters of the framework.
+Because the main content of the prompts is composed of texts from the dataset, and these texts are partitioned in each split, we can assume that the computational cost, in tokens, for this recursion is $O(SD)$, where $S$ is the sample size and $D$ is the maximum tree depth.
+$K$ can be chosen as a regularization parameter avoiding recursing into branches with too few elements, similarly to the regularization of decision trees.
+Importantly, this recursive splitting technique means that, if the trees are balanced, then the maximum depth of the tree is $M_{d}=\lfloor\log_{2}S\rfloor+1$.
+This is an important feature of this technique, as it means that a larger sample allows asking more questions about the same
 Classification
@@
 We then extract the “thematic signature” of this node, defined as the unique sequence of questions and answers along the path from the root.
-This sequence is then translated into a set of structured constraints that are incorporated into a prompt to guide a generative LLM.
+This sequence is then translated into a set of structured constraints that are incorporated into a prompt to guide a generative LLM, leading to a prompt as shown in Figure 5.
+System Prompt You are a skillful writer.
+You are writing under the following context: [context describing the dataset idea, e.g., this is a movie review] Ensure the material is coherent and relevant to the questions asked.
+You will be asked to generate material based on questions and answers Reply with a json string that containing the generated material in the following schema: {schema_str} User Prompt Questions and answers: [list of questions and answers leading to the specific leaf node used for generation]
+Figure 5: Prompt used for CTG. The JSON schema used for prompting helps parsing the answer.
 Experiments and Results
@@
 Importantly, some items in each dataset triggered the LLM’s safety protections due to violent or sexual content. These items were discarded from the sample.
-Qualitative Analysis of Thematic Trees
-We generated thematic trees for each dataset to perform a qualitative analysis of RTP’s behavior [^4] Parameters: sample size 100, $N=4$, $D=5$, $K=3$, no summarization.
-They are shown in figures 2 (IMDB), 3 (Yelp Reviews), and 4 (AG-News).
-We immediately note that the questions seem to be related to the dataset content.
-[root] Does the text primarily provide a positive review of the film or work discussed?
-      [yes] Does the text describe the production or plot of a movie, show, or documentary that involves or depicts a romantic relationship between characters?
-      [yes] Does this text describe a film where the main character faces a personal moral or ethical struggle?
-      [no] Does the text describe a romantic relationship as a central theme in the movie?
-      [yes] Does the text mention a prominent female character in the story?
-      [no] Does the text discuss the cinematography or visual presentation of the film or show?
-      [yes] Does the text mention a documentary or a film about saving the planet or environmental issues?
-      [no] Does the text describe the film in a very positive light overall?
-      [no] Does the text describe a film or show as being emotionally impactful?
-      [yes] Does the text recommend watching the movie or show?
-      [no] Does the text mention a television show or series?
-      [no] Does the text describe the movie as disappointing or poorly executed?
-      [yes] Does the reviewer criticize the script for being poorly written or uninteresting?
-      [yes] Does the text have a positive overall assessment of the movie?
-      [no] Does the text have a positive overall assessment of the movie?
-      [no] Does the text mention disappointment with the film’s acting or storyline?
-      [yes] Does the text criticize the film’s lack of coherence or clear direction?
-      [no] Does the text describe the movie as boring?
-      [no] Does the text mention a train?
-      [no] Does the text mention a specific film director by name?
-      [no] Does the text critique an unjust or exaggerated portrayal of characters or events?
-Figure 2: Thematic tree generated for the IMDB dataset.
-[root] Does the text mention the author having a positive experience with the food or services provided?
-      [yes] Does the review mention being satisfied with the food at the establishment?
-      [yes] Does the text mention a specific dish or food item that the author enjoyed at the establishment?
-      [yes] Does the text mention a specific dish or menu item that the author enjoyed?
-      [yes] Does the text mention a specific dish or menu item that the author enjoyed?
-      [no] Does the text mention positive aspects of the customer service or staff?
-      [yes] Does the reviewer mention exceptional or notable service in the text?
-      [no] Does the text mention a specific issue or drawback related to the described experience or place?
-      [no] Does the text mention dissatisfaction with food or drink quality?
-      [yes] Does the review mention an issue with the food’s preparation or quality?
-      [yes] Does the review mention an issue with the food’s preparation or quality?
-      [no] Does the text describe a place where a purchase can be made?
-      [yes] Does the text mention a specific price for a product or service?
-      [no] Does the text mention experiencing poor customer service?
-      [yes] Does the text describe an experience with poor service?
-      [yes] Does the text describe an experience with poor service?
-      [yes] Does the text describe an experience with poor service?
-      [no] Does the text describe a positive experience with the food at the establishment?
-      [no] Is this review primarily focused on dissatisfaction with the food or beverage quality?
-      [yes] Does the text describe significant dissatisfaction with the food quality?
-      [no] Does the text describe an experience involving food or drink?
-Figure 3: Thematic tree generated for the Yelp Reviews dataset.
-[root] Does the text describe a significant organizational, economic, or structural change?
-      [yes] Does the text mention a plan or agreement to cut jobs, operations, or resources?
-      [no] Does the text mention financial losses or challenges faced by a company?
-      [yes] Does the text mention financial troubles or losses for the company involved?
-      [yes] Does the text report a resolution or agreement in response to a financial or operational challenge?
-      [no] Does the text mention a company cutting, reducing, or altering any of its offerings or practices?
-      [yes] Does the text mention a reduction or change being made to a product or practice?
-      [no] Does the text mention a merger or collaboration between two entities?
-      [no] Does the text mention sports or athletes?
-      [yes] Does the text mention a victory or winning event?
-      [yes] Does the text mention a specific individual achieving victory or recognition?
-      [yes] Does the text mention a victory or win in a sports competition?
-      [no] Does the text describe a decisive event or moment in a sports context?
-      [yes] Does the text specifically mention a score or ranking?
-      [no] Does the text mention an injury?
-      [no] Does the text explicitly mention a financial market or economic activity?
-      [yes] Does the text mention oil prices or crude oil?
-      [yes] Does the text mention oil prices changing?
-      [no] Does the text mention a financial or economic statistic or figure?
-      [no] Does the text mention a government action or decision?
-      [yes] Does the text mention an instance of a government or official body negotiating or responding to a conflict or dispute?
-      [no] Does the text mention a specific event occurring in nature or the environment?
-Figure 4: Thematic tree generated for the AG-News dataset.
-A qualitative examination of the trees generated by RTP reveals several key characteristics of our method, highlighting both its strengths in thematic discovery and its interesting emergent behaviors.
+Hyperparameters
+In these experiments, we used the following hyperparameters:
+  • Sample size $S=100$, which we found to be high enough for a reasonable balanced tree depth $M_{d}=7$, and, simultaneously, allowing all elements withing the sample to fit within the LLM’s context window;
+  • Maximum depth $D=5$, which is less than $M_{d}$ and aims to produce nodes with an average of $4$ samples;
+  • Minimum elements per node $K=3$, avoiding leaf nodes that are too specific and become non-representative.
+  • Number of answers $N=4$, aiming to stabilize answers without significantly increasing the token count.
+These hyperparameters were chosen to favor particular sizes for the trees. A full empirical study of their effect in the tree induction process is a key direction for future work.
+Analysis of Thematic Trees
+We generated thematic trees for each dataset to perform a qualitative analysis of RTP’s behavior.
+The token usage is shown in Table 1.
+As we can see, token usage for the experiments was in the range of $10^{5}$ to $10^{6}$, hence costing less than 1 USD with current API princing.
+Dataset Q A Total
+IMDB 135204 231036 366240
+Yelp 94629 192993 287622
+AGNews 29615 122667 152282
+Table 1: Token usage for each experiment in the question creation stage (Q), the question answering stage (A), and total.
+They are shown in figures 6 (IMDB), 7 (Yelp Reviews), and 8 (AG-News) (located in the appendix). We immediately note that the questions seem to be related to the dataset content.
+Alignment with ground truth labels
+To quantitatively analyze the alignment between RTP’s unsupervised partitions and the ground-truth labels, we employ two complementary metrics for each node: Information Entropy and Node Purity.
+First, we use Information Entropy, a standard measure of impurity from information theory.
+A low entropy value indicates that a node contains documents from a limited number of classes, with a value of 0 representing a perfectly ”pure” node.
+It is calculated as:
+$H(S)=-\sum_{i=1}^{c}p_{i}\log_{2}(p_{i})$ (1)
+where $S$ is the set of documents in a given node, $c$ is the number of unique classes, and $p_{i}$ is the proportion of documents in $S$ belonging to class $c_{i}$.
+Second, to provide a more direct and intuitive measure, we calculate Node Purity as the proportion of the majority class within the node:
+$\text{Purity}(S)=\max_{i}(p_{i})$ (2)
+While entropy captures the overall class distribution, purity offers a straightforward assessment of the dominance of a single class.
+Together, these metrics allow us to quantify the emergent correlation between the semantic themes discovered by the LLM and the latent class structure of the data.
+For instance, a high purity and low entropy in a leaf node indicate that the path of questions leading to it has successfully isolated a thematically coherent group that also shares a class label.
+Since RTP is an unsupervised framework that receives no label information during tree construction, perfect purity in all leaves is not the expected outcome.
+Instead, these metrics provide a quantitative lens that complements the following qualitative examination of the trees, which reveals the specific nature of these thematic discoveries and other emergent behaviors of our method.
 Hierarchical Theme Discovery
@@
 We also noted instances of similar or repeated questions appearing in the tree, either in parent-child relationships or between sibling nodes.
+The redundant questions seem to appear in nodes whose purity is higher, which indicates this is linked to biases found within the nodes themselves.
 This is a natural consequence of the algorithm’s design, where question generation for each node is an independent process that only considers the local document partition.
@@
 To provide a fair comparison with a standard topic model, we executed BERTopic in the same dataset sample used in the experiments with RTP.
-The results are shown in Tables 1, 2, and 3.
+The results are shown in Tables 5, 6, and 7, provided in the Appendix.
 Following BERTopic convention, topics are numbered from 0 to the number of topics found, and topic 1-1 represents outliers.
-Topic Representation
--1 br movie like just really film best actors time joseph
-0 br film movie like just bad good really movies thing
-1 film br movie police hes mukhsin tolstoy way good love
-2 film musical best brando singing dancing george burns astaire fontaine
-3 br action point dont sasha film does believe like money
-4 ethel movie dead story advancement backwoods bodies battalion halfway film
-5 color life natali films think movie important based white kind
-Table 1: Topics generated by BERTopic in IMDB
-Topic Representation
--1 food like really just good place groupon ordered lunch told
-0 bar got airport way people air like line view terminal
-1 great good delicious ordered place places crepes breakfast lunch phoenix
-2 food thai good great pho place menu hair love felt
-3 room told hotel desk card security stay roland didnt safe
-4 table horrible minutes busy servers sat right server service del
-5 pizzas pizza papa called crust toppings johns driver looked time
-6 wash sushi mat creme includes vacuum price machine priced tea
-7 salmon hype fruit place garlic dont flavors instead fries time
-8 salad cardboard hickory food salads piece know dressing husband came
-9 buffet wait quality seated relleno chile youre enchiladas beans food
-10 sandwiches night cute owner nni coffee lots quite huge big
-Table 2: Topics generated by BERTopic in Yelp Reviews
-Topic Representation
--1 ap monday bush biggest said solaris group british hp 39s
-0 39s reuters oil corp prices cut stocks said new york
-1 election iran nuclear said afp north ltbgtltbgt syria sudan region
-2 lions 39t inter williams win sunday season standings dallas revolution
-3 yankees sox look jeter games johnson peskys rays aftermath charleys
-4 39s england world bowler open number test coach champions harmison
-5 bison bears falls taj river salmon mahal mcneil spectacular sun
-6 alitalia alter emissions end split practices canada unions deal sales
-Table 3: Topics generated by BERTopic in AG-News
 Analysis
@@
 This ensures a fair comparison, isolating the benefit of RTP’s structure against the fine-tuning of a powerful neural network on the same limited data.
-For context, we also report the state-of-the-art (SoTA) performance from models on huggingface.com  [^5] IMDB: https://huggingface.co/DeepAxion/distilbert-imdb-sentiment; Yelp Reviews: https://huggingface.co/BexRedpill/distilbert-on-yelp-reviews-full-epoch-2; AG-News: https://huggingface.co/JiaqiLee/bert-agnews, which are typically trained on the full dataset.
+For context, we also report the state-of-the-art (SoTA) performance among all models available on huggingface.com  [^4] We manually sought models fine-tuned in the target datasets IMDB: https://huggingface.co/DeepAxion/distilbert-imdb-sentiment; Yelp Reviews: https://huggingface.co/BexRedpill/distilbert-on-yelp-reviews-full-epoch-2; AG-News: https://huggingface.co/JiaqiLee/bert-agnews.
+These results are compatible with those found in recent Kaggle competition leaderboards like https://www.kaggle.com/competitions/sentiment-analysis-on-imdb-dataset-2025/leaderboard and https://www.kaggle.com/competitions/enap-bert-para-classificao/leaderboard., which are typically BERT-based transformers (distillBERT for IMDB and Yelp, BERT for AG-News) trained on the full dataset.
+We include these scores not as a direct competitive benchmark, as our protocol is intentionally data-scarce, but as a reference point to contextualize our results.
 Results
-Our primary finding is that RTP significantly outperforms the fine-tuned DistilBERT baseline in IMDB, and surpasses the performance of a model trained on the full dataset ($p<0.05$, one-sample t-test).
+Our primary finding is that RTP significantly outperforms the fine-tuned DistilBERT baseline in IMDB.
 The performances of RTP and the baseline are similar in Yelp Reviews.
 However, on the topic-based AG News dataset, the fine-tuned DistilBERT baseline is superior.
-This brings forward a key characteristic of our method.
 Model IMDB Yelp Reviews AG-News
 SoTA (reference) .94 .65 .94
-DistillBERT (baseline) .81(.03) .39(.07) 0.87(.01)
+DistillBERT (baseline) .81(.03) .39(.07) .87(.01)
 RTP (ours) 96(.02) .40(.12) .64(.04)
@@
 The highest mean accuracy (excluding SoTA) is marked in bold.
-Values are $\mu(2\sigma)$.
-Asterisks denote non-significant differences ($p>.05$) compared to baseline.
+Values are $\mu(c)$, where $c$ is the $95\%$ confidence interval.
+Asterisks denote non-significant differences ($p>.05$ using a t-test) compared to baseline.
 Analysis
@@
 Exploring interactive interfaces that allow users to guide the tree-building process could further enhance the model’s alignment with specific analytical goals, bridging the gap between automated discovery and expert-driven knowledge construction.
+Appendix A Appendix: Thematic trees for all executions
+Figure 6: Thematic tree generated for the IMDB dataset.
+Figure 7: Thematic tree generated for the Yelp Reviews dataset.
+Figure 8: Thematic tree generated for the AG-News dataset.
+Appendix B Appendix: Topics generated by BERTopic
+Topic Representation
+-1 br movie like just really film best actors time joseph
+0 br film movie like just bad good really movies thing
+1 film br movie police hes mukhsin tolstoy way good love
+2 film musical best brando singing dancing george burns astaire fontaine
+3 br action point dont sasha film does believe like money
+4 ethel movie dead story advancement backwoods bodies battalion halfway film
+5 color life natali films think movie important based white kind
+Table 5: Topics generated by BERTopic in IMDB
+Topic Representation
+-1 food like really just good place groupon ordered lunch told
+0 bar got airport way people air like line view terminal
+1 great good delicious ordered place places crepes breakfast lunch phoenix
+2 food thai good great pho place menu hair love felt
+3 room told hotel desk card security stay roland didnt safe
+4 table horrible minutes busy servers sat right server service del
+5 pizzas pizza papa called crust toppings johns driver looked time
+6 wash sushi mat creme includes vacuum price machine priced tea
+7 salmon hype fruit place garlic dont flavors instead fries time
+8 salad cardboard hickory food salads piece know dressing husband came
+9 buffet wait quality seated relleno chile youre enchiladas beans food
+10 sandwiches night cute owner nni coffee lots quite huge big
+Table 6: Topics generated by BERTopic in Yelp Reviews
+Topic Representation
+-1 ap monday bush biggest said solaris group british hp 39s
+0 39s reuters oil corp prices cut stocks said new york
+1 election iran nuclear said afp north ltbgtltbgt syria sudan region
+2 lions 39t inter williams win sunday season standings dallas revolution
+3 yankees sox look jeter games johnson peskys rays aftermath charleys
+4 39s england world bowler open number test coach champions harmison
+5 bison bears falls taj river salmon mahal mcneil spectacular sun
+6 alitalia alter emissions end split practices canada unions deal sales
+Table 7: Topics generated by BERTopic in AG-News